OpenStudy (anonymous):

Can someone please tell me how to find the area of one slice of pizza if my pizza is 12” in diameter and has 12 slices?

hartnn (hartnn):

can u find area of circle when diameter is given ?

OpenStudy (anonymous):

I just don't know what to do, no.

hartnn (hartnn):

can u find radius when diameter is given ?

hartnn (hartnn):

no ?

OpenStudy (anonymous):

No? Can you just show me the steps?

OpenStudy (anonymous):

I know you do: Interior angle = 360°/12 = 30° Arc length = a/360°(πd)

OpenStudy (anonymous):

but then i'm stuck

hartnn (hartnn):

yes, radius = diameter /2 = 12/2 = 6

hartnn (hartnn):

first we can find area of circle, then divide it by 12

hartnn (hartnn):

area of circle = pi * r^2

hartnn (hartnn):

area of slice = pi * r^2/12 r= 6

OpenStudy (anonymous):

i'm confused, okay. i'm sorry, i need to find the interior angle and arc length of the slices AND the area of each slice. if my pizza is 12" in diameter and has 12 slices.

OpenStudy (anonymous):

Area = π(r)2 so pizza area = 3.14*6*6 = 113.04 inch sq for 12 the part = 113.04/12 = 9.42 sq in

OpenStudy (anonymous):

piz·za [ptsə] (plural piz·zas) noun baked dough topped with cheese: a flat round piece of bread dough baked with a variety of toppings, often including tomato sauce and cheese [Late 19th century. < Italian, "pie"]

OpenStudy (anonymous):

Okay so 9.42 sq inch per slice but now i need to find the interior angle and arc length of each slice..

hartnn (hartnn):

arc length of circle = ? then just divide it by 12 your interior angle = 30 is correct

hartnn (hartnn):

so arc length of slice = r*360/12 r= radius

OpenStudy (anonymous):

the angle will be 360/12 = 30 degree the arc length also easy as \[2 \pi r\] = 2*3.14*6 = 37.68 inch so for this 12th part will be 37.68/12 = 3.14 inch only

OpenStudy (anonymous):

okay i was doing this: interior angle = 360°/12 = 30° arc length = 30°/360° (π • 12) arc length = 0.0833333333333333 (π • 12) can you tell me what i was doing wrong or what i need to do ahhh

OpenStudy (anonymous):

see above then u can understand or do some drawing in rough book

OpenStudy (anonymous):

i did look above and it just makes me confused, can you please tell me straight out what i have to do left? i'm sorry :(

hartnn (hartnn):

u did nothing wrong, arc length = 0.0833333333333333 (π • 12)=3.1414 inch which is correct

hartnn (hartnn):

got this? from arc length , u want area ?

OpenStudy (anonymous):

yeah what's the next step?

OpenStudy (anonymous):

ohh i see i see

hartnn (hartnn):

to get area from arc length , u multiply it by (r/2) here r=6

OpenStudy (anonymous):

So this is interior angle and arc length of each slice: interior angle = 360°/12 = 30° arc length = 30°/360° (π • 12) arc length = 0.0833333333333333 (π • 12) = 3.1414 inch

hartnn (hartnn):

yes..

OpenStudy (anonymous):

thank you so much oh gosh you've been a great help really, both of you

hartnn (hartnn):

welcome ^_^